Find 3 consecutive odd integers such that the smallest is 10 more than the sum of the other two integers.
Answer by longjonsilver(2297)

let smallest integer be x
Middle one is therefore x+2
and largest is x+4
smallest = 10 + (other 2 added)
--> x = 10 + (x+2) + (x+4)
x = 16 + 2x
-x = 16
x = -16
so numbers are -16, -14 and -12.

However, these are not odd, so i would check the wording in your question.

jon